If the equation of a parabola with the focus at (3,−4) and the directrix x+y = 2 is
x
2 +y
2 −2xy−8x+20y+c = 0, then what is the value of c
1
Expert's answer
2020-03-02T10:31:20-0500

Concept:


For a parabola, axis of parabola is perpendicular to the directrix. we also know distance from parabola to its focus and directix are equal which means DV= VF i.e. V is the mid point of the perpendicular line connecting focus F and to the directrix at D .

--------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Let focus (F) = (3, -4)

Since equation of directrix is known

"x+y=2"

or "y= -x +2"

slope of directrix = -1

since for two perpendicular line having slope m1 and m2, "m_1*m_2=-1"

slope of axis of parabola = -1/(slope of directrix) =-1/(-1) = 1

equation of axis of parabola will be line having slope =1 and passing through focus F=(3, -4)

"y-(-4)=1(x-3)"

"y=x-7"


Axis will met the directrix at point D which coordinate will be found by solving equation "y=-x+2" and "y=x-7"

adding two equation

"2y = -5"

"y=-5\/2"

substituting value in first equation

"-5\/2=-x+2"

"x=9\/2"

Thus point D = ( 9/2 , -5/2)

Coordinate of point V using mid point formula

"V=(\frac{3+9/2}{2}, \frac{-4-5/2}{2})=(\frac{15}{4}, -\frac{13}{4})"

Since V lies on parabola, it will satisfy the equation of parabola

"x^2+y^2-2xy-8x+20y+c=0"

"\left(\frac{15}{4}\right)^2+\left(-\frac{13}{4}\right)^2-2\left(\frac{15}{4}\right)\left(-\frac{13}{4}\right)-8\left(\frac{15}{4}\right)+20\left(-\frac{13}{4}\right)+c=0"

"\frac{225}{16}+\frac{169}{16}+\frac{195}{8}-30-65+c=0"

multiply both side by 16

"225+169+390-480-1040+16c=0"

"-736+16c=0"

"16c=736"

"c=736\/16 = 46"

c = 46 is the answer.